OpenStudy (anonymous):

The roots of equation 5^2x + (-30)5^x + 125 =0 ?

OpenStudy (anonymous):

let 5^x = X then X^2 - 30X + 125 = 0 (X -5)(X - 25) = 0 X = 5 and X = 25 5^x = 5 , 25 find x

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

x= 1 , x=2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thxx a lot

OpenStudy (anonymous):

glad to be of help